# The Prism Hypothesis: Harmonizing Semantic and Pixel Representations via Unified Autoencoding

### v1

@@ 1 Introduction
Trained on massive corpora, recent foundation models have profoundly reshaped perception and generation systems, generalizing well across diverse downstream tasks [dosovitskiy2021vit, caron2021dino, oquab2023dinov2, radford2021clip]. Yet, early advances in perception and generation evolve along largely separate trajectories. Their objectives are typically distributed across distinct network structures, e.g., employing pretrained semantic encoders [caron2021dino, oquab2023dinov2, radford2021clip], to capture high-level meaning, or pixel encoders [kingma2014vae, razavi2019vqvae2] to compress fine-grained visual detail. While each module excels within its own domain, this fragmentation compels subsequent unification efforts [qu2025tokenflow, huang2025illume+, tang2025unilip] to depend simultaneously on semantic and pixel encoders, forcing networks to reconcile fundamentally heterogeneous representations. The sharp mismatch lowers training efficiency and induces representational conflicts, with these incompatible features often interfering rather than complementing one another.
This fragmentation uncovers a deep-seated tension between abstraction and fidelity, which is a driving force in shaping subsequent foundation models. To alleviate it, recent studies [zheng2025rae, yue2025uniflow, shi2025svg] attempt to transfer semantic encoders into visual generation domains alongside strong pixel decoders. This strategy substantially accelerates convergence and improves semantic correspondence, yet remains limited in recovering fine-grained visual details. In parallel, another research direction seeks to endow pixel encoders with semantic awareness through text supervision [xu2022groupvit, yi2023simseg, wu2024imagetext], semantic encoder distillation [chen2023zeroseg, huang2023g2sd], and hierarchical feature integration [wu2024imagetext]. While these efforts move toward unifying understanding and generation representations in a single module, they often achieve coexistence through trade-offs rather than genuine integration.
@@
From this perspective, we introduce Unified Autoencoding (UAE), a tokenizer that learns a shared latent space and harmonizes semantic structure and pixel-level detail. Specifically, UAE features an innovative frequency-band modulator that factorizes real-world content into a fundamental semantic band and residual pixel bands with controllable fine granularity. This design is not only grounded in comprehensive empirical evidence but also validated across diverse reconstruction and perception tasks. With its compact yet semantically expressive representations, UAE outperforms concurrent RAE [zheng2025rae], SVG [shi2025svg], and UniFlow [yue2025uniflow] across rFID, PSNR, gFID and Accuracy metrics on ImageNet, demonstrating that its learned latent space is both semantically representative and pixel-faithful.
@@ 2 Related Work
Unifying the representations between pixel and semantic embeddings has become a central objective for existing foundation models. Joint embedding approaches align images and text in a shared representation and enable strong zero-shot transfer [radford2021clip, jia2021align], and have been extended to many modalities, e.g., audio, depth, thermal, and inertial signals [girdhar2023imagebind]. In parallel, modality-agnostic backbones aim to build a unified architecture that can process diverse input modalities and generate task-specific outputs through learned queries or a shared token representation [jaegle2021perceiverio, lu2022unifiedio, lu2024unifiedio2, wang2023beit3, baevski2022data2vec].
On the tokenizer side, discrete codebook methods have demonstrated that the design and granularity of visual tokens play a crucial role in determining how effectively a single sequence model can adapt to vision tasks [esser2021taming, yu2024magvitv2]. Recent work goes further and seeks tokenizers that support both understanding and generation at the same time. OmniTokenizer [wang2024omnitokenizer] learns a joint image video tokenizer with a spatial-temporal decoupled transformer and reports strong reconstruction and synthesis across both domains. Very recent studies deepen this trend. UniFlow [yue2025uniflow] proposes a unified pixel flow tokenizer that adapts a pretrained encoder through layer-wise self-distillation and employs a lightweight patch-wise flow decoder, explicitly targeting the long-standing tension between semantic abstraction and pixel-faithful reconstruction. Two concurrent works remove the traditional variational bottleneck entirely and build unified representation latents for diffusion transformers. Diffusion Transformers with Representation Autoencoders (RAE) [zheng2025rae] replace the usual reconstruction-only encoder with a pretrained representation encoder, e.g., DINO or SigLIP, and a trained decoder, arguing that semantically rich latents accelerate convergence and improve generative fidelity. SVG [shi2025svg] trains a diffusion model on DINO features with a small residual branch for details, reporting faster training, few-step sampling, and improved quality.
@@
Classical image synthesis adopted pyramids and wavelets to separate structure by scale, enabling coarse-to-fine generation and targeted refinement of detail. A typical example is the Laplacian pyramid of adversarial networks [denton2015lapgan], which trains a generator per level and synthesizes images by successively adding higher-frequency residuals. Subsequent analyses of neural networks from a spectral perspective showed that standard architectures prioritize low frequencies and learn higher frequencies later, a phenomenon known as spectral bias [rahaman2019spectral]. Two lines of work respond to this bias. The first uses input or architecture design to improve access to high frequencies, e.g., Fourier feature mappings and periodic activations that help multi-layer perceptrons represent fine detail [tancik2020fourier, sitzmann2020siren]. The second introduces frequency-aware objectives and signal processing choices, such as focal frequency loss to emphasize hard frequencies and alias-free synthesis to avoid spurious high-frequency artifacts [jiang2021ffl, karras2021aliasfree].
Modern generative models retain this multi-resolution view. Cascaded diffusion trains models at increasing resolutions, allowing each stage to learn the appropriate frequency band and its own error distribution [ho2022cascaded]. Variants construct explicit feature pyramids or hierarchical patch schedules, enabling efficient diffusion on large images and video while preserving high-frequency detail [fan2023frido, skorokhodov2024hpdm]. Recent latent diffusion designs introduce cross-magnification spaces or zoomable pyramids that share information across scales and enable large-image reconstruction without retraining [yellapragada2025zoomldm]. In autoregressive models, VAR [tian2024var] casts generation as predicting the next scale or resolution, showing strong ImageNet results and clean scaling trends. This explicit progression from global layout to fine detail emerges as a viable alternative to diffusion. Building on this idea, Next Visual Granularity (NVG) generation [wang2025nvg] produces sequences at a fixed resolution but with progressively finer token granularity, surpassing prior VAR baselines while maintaining structured control over detail. Similarly, NFIG [huang2025nfig] performs discrete next-frequency prediction, demonstrating strong generative performance on the ImageNet benchmark. Our approach aligns with these trends but focuses on unified representation rather than the generator’s training schedule.
@@ 3.3 Residual Split Flow
Given a latent grid $\mathbf{z}\in\mathbb{R}^{B\times C\times H\times W}$, we decompose $\mathbf{z}$ into $K$ frequency bands $\mathbf{z_{f}}\in\mathbb{R}^{B\times K\times C\times H\times W}$ using an FFT Band Projector followed by an Iterative Split procedure. Inspired by the coupling mechanism in flow-based models such as Glow [kingma2018glow], the decomposition is performed in the frequency domain via FFT-based projection, allowing each band to capture a distinct spectral component while maintaining invertibility and spatial consistency.
@@ 4.1 Implementation Details
Training Settings. We train UAE with DINOv2-B and DINOv2-L respectively in 256x256 resolution. The training processes are decomposed into three stages. In stage 1, the semantic encoder is frozen, where the decoder is trained by reconstruction loss. In stage 2, we unfreeze the encoder and finetune the full model with semantic-wise loss and reconstruction loss. In stage 3, we finetune the full model end-to-end with noise injection and GAN loss, where the GAN loss setting follows the representative RAE [zheng2025rae].
Table 1: Reconstruction quality on ImageNet-1K and MS-COCO 2017 (256x256). Note that our proposed UAE achieves state-of-the-art reconstruction among unified tokenizers and remains competitive with strong autoencoders such as Flux-VAE and SD3-VAE. Under identical DINOv2 encoders, UAE significantly outperforms the RAE baseline in both PSNR and SSIM while reducing rFID by over 90%. When scaled to DINOv2-L, UAE attains the best overall perceptual quality (rFID=0.16) and fidelity (PSNR=33.08, SSIM=0.94), demonstrating the effectiveness of frequency-aware factorization in preserving both semantic and fine-grained visual detail.
@@
SD-VAE [rombach2021highresolution] Continuous 8 25.68 0.72 0.75 25.43 0.73 0.76
SD-VAE-EMA [rombach2021highresolution] Continuous 8 24.99 0.71 0.63 24.76 0.72 0.57
SD3-VAE [lopez2025sd3] Continuous 8 29.58 0.86 0.21 29.50 0.87 0.19
Flux-VAE [labs2025flux1kontextflowmatching] Continuous 8 31.07 0.89 0.17 30.99 0.90 0.19
TokenFlow [qu2025tokenflow] Discrete 16 21.41 0.69 1.37 - - -
DualViTok [huang2025illume+] Discrete 16 22.53 0.74 1.37 - - -
EMU2 [Emu2] Continuous 14 13.49 0.42 3.27 - - -
UniLIP [tang2025unilip] Continuous 32 22.99 0.75 0.79 - - -
@@ 4.2 Visual Reconstruction
Quantitative Evaluation. In Table 1, We quantify reconstruction quality at $256\times 256$ on ImageNet-1K and MS-COCO 2017. The compared baselines include widely-used generative tokenizers and variational decoders. For a fair comparison, we report the evaluation results for our UAE and the corresponding baseline that follows the same configurations of DINOv2-base. Here, we report PSNR and SSIM for fidelity, and rFID for perceptual quality.
Figure 5: Qualitative comparison of reconstruction fidelity across autoencoding paradigms. We visualize reconstructed samples from representative methods, including SD-VAE [rombach2022ldm], RAE [zheng2025rae], and our proposed UAE. Each row corresponds to reconstructions from a fixed source set spanning text, human, object, and artistic domains. UAE produces the most consistent and semantically faithful reconstructions, preserving both high-frequency details (e.g., texture and edge sharpness) and global structure (e.g., layout and color harmony), while reducing the blurring and semantic drift observed in SD-VAE and RAE. (The detail comparisons are denoted in the yellow boxes.)
@@ 4.3 Generative Modeling
DiT-XL/2 [peebles2023dit] 2.27 271.3 0.83 0.57
VAR [tian2024var] 1.73 350.2 0.82 0.6
UniFlow [yue2025uniflow] 2.45 228.0 - -
RAE [zheng2025rae] 1.51 242.9 0.79 0.63
@@
We conduct all generative experiments in the multi-band latent space. All experimental settings follow those used in RAE [zheng2025rae] to ensure a fair comparison. As shown in Table 2, our UAE attains a gFID of 1.68 and an IS of 301.6, achieving performance on par with existing state-of-the-art generative models. This suggests that the unified frequency-based representation enables the generative model to progressively capture both global structure and fine-grained details in a coherent manner, keeping highly generative while preserving strong semantic quality. Overall, the UAE latent space, constructed through explicit low- and high-frequency decomposition, provides an effective and diffusion-friendly foundation for large-scale visual generation.
@@ 4.4 Semantic Understanding
To evaluate the learned semantic discriminability, we perform linear probing on ImageNet-1K following the standard protocol in [yue2025uniflow]. All encoder weights are frozen, and a single linear classifier is trained on top of the global embeddings for 30 epochs using SGD with a cosine learning rate schedule and a base rate of 0.1. Consistent augmentations are applied across methods for a fair comparison.
@@
VFMTok [zheng2025vision] ViT-L 69.4
BEiT [beit] ViT-L 73.5
MAE [he2022mae] ViT-L 75.8
MAGE [li2023mage] ViT-L 78.9
UniFlow [yue2025uniflow] ViT-L 82.6
RAE [zheng2025rae] ViT-B 83.0

### v2

@@ 1 Introduction
Trained on massive corpora, recent foundation models have profoundly reshaped perception and generation systems, generalizing well across diverse downstream tasks [25, 3, 26, 29]. Yet, early advances in perception and generation evolve along largely separate trajectories. Their objectives are typically distributed across distinct network structures, e.g., employing pretrained semantic encoders [3, 26, 29], to capture high-level meaning, or pixel encoders [18, 31] to compress fine-grained visual detail. While each module excels within its own domain, this fragmentation compels subsequent unification efforts [28, 11, 39] to depend simultaneously on semantic and pixel encoders, forcing networks to reconcile fundamentally heterogeneous representations. The sharp mismatch lowers training efficiency and induces representational conflicts, with these incompatible features often interfering rather than complementing one another.
This fragmentation uncovers a deep-seated tension between abstraction and fidelity, which is a driving force in shaping subsequent foundation models. To alleviate it, recent studies [50, 48, 34] attempt to transfer semantic encoders into visual generation domains alongside strong pixel decoders. This strategy substantially accelerates convergence and improves semantic correspondence, yet remains limited in recovering fine-grained visual details. In parallel, another research direction seeks to endow pixel encoders with semantic awareness through text supervision [44, 46, 43], semantic encoder distillation [4, 12], and hierarchical feature integration [43]. While these efforts move toward unifying understanding and generation representations in a single module, they often achieve coexistence through trade-offs rather than genuine integration.
@@
From this perspective, we introduce Unified Autoencoding (UAE), a tokenizer that learns a shared latent space and harmonizes semantic structure and pixel-level detail. Specifically, UAE features an innovative frequency-band modulator that factorizes real-world content into a fundamental semantic band and residual pixel bands with controllable fine granularity. This design is not only grounded in comprehensive empirical evidence but also validated across diverse reconstruction and perception tasks. With its compact yet semantically expressive representations, UAE outperforms concurrent RAE [50], SVG [34], and UniFlow [48] across rFID, PSNR, gFID and Accuracy metrics on ImageNet, demonstrating that its learned latent space is both semantically representative and pixel-faithful.
@@ 2 Related Work
Unifying the representations between pixel and semantic embeddings has become a central objective for existing foundation models. Joint embedding approaches align images and text in a shared representation and enable strong zero-shot transfer [29, 15], and have been extended to many modalities, e.g., audio, depth, thermal, and inertial signals [8]. In parallel, modality-agnostic backbones aim to build a unified architecture that can process diverse input modalities and generate task-specific outputs through learned queries or a shared token representation [14, 24, 23, 42, 1].
On the tokenizer side, discrete codebook methods have demonstrated that the design and granularity of visual tokens play a crucial role in determining how effectively a single sequence model can adapt to vision tasks [6, 47]. Recent work goes further and seeks tokenizers that support both understanding and generation at the same time. OmniTokenizer [41] learns a joint image video tokenizer with a spatial-temporal decoupled transformer and reports strong reconstruction and synthesis across both domains. Very recent studies deepen this trend. UniFlow [48] proposes a unified pixel flow tokenizer that adapts a pretrained encoder through layer-wise self-distillation and employs a lightweight patch-wise flow decoder, explicitly targeting the long-standing tension between semantic abstraction and pixel-faithful reconstruction. Two concurrent works remove the traditional variational bottleneck entirely and build unified representation latents for diffusion transformers. Diffusion Transformers with Representation Autoencoders (RAE) [50] replace the usual reconstruction-only encoder with a pretrained representation encoder, e.g., DINO or SigLIP, and a trained decoder, arguing that semantically rich latents accelerate convergence and improve generative fidelity. SVG [34] trains a diffusion model on DINO features with a small residual branch for details, reporting faster training, few-step sampling, and improved quality.
@@
Classical image synthesis adopted pyramids and wavelets to separate structure by scale, enabling coarse-to-fine generation and targeted refinement of detail. A typical example is the Laplacian pyramid of adversarial networks [5], which trains a generator per level and synthesizes images by successively adding higher-frequency residuals. Subsequent analyses of neural networks from a spectral perspective showed that standard architectures prioritize low frequencies and learn higher frequencies later, a phenomenon known as spectral bias [30]. Two lines of work respond to this bias. The first uses input or architecture design to improve access to high frequencies, e.g., Fourier feature mappings and periodic activations that help multi-layer perceptrons represent fine detail [38, 35]. The second introduces frequency-aware objectives and signal processing choices, such as focal frequency loss to emphasize hard frequencies and alias-free synthesis to avoid spurious high-frequency artifacts [16, 17].
Modern generative models retain this multi-resolution view. Cascaded diffusion trains models at increasing resolutions, allowing each stage to learn the appropriate frequency band and its own error distribution [10]. Variants construct explicit feature pyramids or hierarchical patch schedules, enabling efficient diffusion on large images and video while preserving high-frequency detail [7, 36]. Recent latent diffusion designs introduce cross-magnification spaces or zoomable pyramids that share information across scales and enable large-image reconstruction without retraining [45]. In autoregressive models, VAR [40] casts generation as predicting the next scale or resolution, showing strong ImageNet results and clean scaling trends. This explicit progression from global layout to fine detail emerges as a viable alternative to diffusion. Building on this idea, Next Visual Granularity (NVG) generation [wang2025nvg] produces sequences at a fixed resolution but with progressively finer token granularity, surpassing prior VAR baselines while maintaining structured control over detail. Similarly, NFIG [13] performs discrete next-frequency prediction, demonstrating strong generative performance on the ImageNet benchmark. Our approach aligns with these trends but focuses on unified representation rather than the generator’s training schedule.
@@ 3.3 Residual Split Flow
Given a latent grid $\mathbf{z}\in\mathbb{R}^{B\times C\times H\times W}$, we decompose $\mathbf{z}$ into $K$ frequency bands $\mathbf{z_{f}}\in\mathbb{R}^{B\times K\times C\times H\times W}$ using an FFT Band Projector followed by an Iterative Split procedure. Inspired by the coupling mechanism in flow-based models such as Glow [19], the decomposition is performed in the frequency domain via FFT-based projection, allowing each band to capture a distinct spectral component while maintaining invertibility and spatial consistency.
@@ 4.1 Implementation Details
Training Settings. We train UAE with DINOv2-B and DINOv2-L respectively in 256x256 resolution. The training processes are decomposed into three stages. In stage 1, the semantic encoder is frozen, where the decoder is trained by reconstruction loss. In stage 2, we unfreeze the encoder and finetune the full model with semantic-wise loss and reconstruction loss. In stage 3, we finetune the full model end-to-end with noise injection and GAN loss, where the GAN loss setting follows the representative RAE [50].
Table 1: Reconstruction quality on ImageNet-1K and MS-COCO 2017 full set (256x256). Note that our proposed UAE achieves state-of-the-art reconstruction among unified tokenizers and remains competitive with strong autoencoders such as Flux-VAE and SD3-VAE. Under identical DINOv2 encoders, UAE significantly outperforms the RAE baseline in both PSNR and SSIM while reducing rFID by over 90%. When scaled to DINOv2-L, UAE attains the best overall perceptual quality (rFID=0.16) and fidelity (PSNR=33.08, SSIM=0.94), demonstrating the effectiveness of frequency-aware factorization in preserving both semantic and fine-grained visual detail.
@@
SD-VAE [32] Continuous 8 25.68 0.72 0.75 25.43 0.73 0.76
SD-VAE-EMA [32] Continuous 8 24.99 0.71 0.63 24.76 0.72 0.57
SD3-VAE [22] Continuous 8 29.58 0.86 0.21 29.50 0.87 0.19
Flux-VAE [20] Continuous 8 31.07 0.89 0.17 30.99 0.90 0.19
TokenFlow [28] Discrete 16 21.41 0.69 1.37 - - -
DualViTok [11] Discrete 16 22.53 0.74 1.37 - - -
EMU2 [37] Continuous 14 13.49 0.42 3.27 - - -
UniLIP [39] Continuous 32 22.99 0.75 0.79 - - -
@@ 4.2 Visual Reconstruction
Quantitative Evaluation. In Section 4.1, We quantify reconstruction quality at $256\times 256$ on ImageNet-1K and MS-COCO 2017. The compared baselines include widely-used generative tokenizers and variational decoders. For a fair comparison, we report the evaluation results for our UAE and the corresponding baseline that follows the same configurations of DINOv2-base. Here, we report PSNR and SSIM for fidelity, and rFID for perceptual quality.
Figure 5: Qualitative comparison of reconstruction fidelity across autoencoding paradigms. We visualize reconstructed samples from representative methods, including SD-VAE [33], RAE [50], and our proposed UAE. Each row corresponds to reconstructions from a fixed source set spanning text, human, object, and artistic domains. UAE produces the most consistent and semantically faithful reconstructions, preserving both high-frequency details (e.g., texture and edge sharpness) and global structure (e.g., layout and color harmony), while reducing the blurring and semantic drift observed in SD-VAE and RAE. (The detail comparisons are denoted in the yellow boxes.)
@@ 4.3 Generative Modeling
DiT-XL/2 [27] 2.27 271.3 0.83 0.57
VAR [40] 1.73 350.2 0.82 0.6
UniFlow [48] 2.45 228.0 - -
RAE [50] 1.51 242.9 0.79 0.63
@@
We conduct all generative experiments in the multi-band latent space. All experimental settings follow those used in RAE [50] to ensure a fair comparison. As shown in Table 2, our UAE attains a gFID of 1.68 and an IS of 301.6, achieving performance on par with existing state-of-the-art generative models. This suggests that the unified frequency-based representation enables the generative model to progressively capture both global structure and fine-grained details in a coherent manner, keeping highly generative while preserving strong semantic quality. Overall, the UAE latent space, constructed through explicit low- and high-frequency decomposition, provides an effective and diffusion-friendly foundation for large-scale visual generation.
@@ 4.4 Semantic Understanding
To evaluate the learned semantic discriminability, we perform linear probing on ImageNet-1K following the standard protocol in [48]. All encoder weights are frozen, and a single linear classifier is trained on top of the global embeddings for 30 epochs using SGD with a cosine learning rate schedule and a base rate of 0.1. Consistent augmentations are applied across methods for a fair comparison.
@@
VFMTok [49] ViT-L 69.4
BEiT [2] ViT-L 73.5
MAE [9] ViT-L 75.8
MAGE [21] ViT-L 78.9
UniFlow [48] ViT-L 82.6
RAE [50] ViT-B 83.0